Several intelligence tests follow a normal distribution with a mean of 100 and a standard deviation of 15.

a) Determine the percentage of the population that would obtain a score between 95 and 110.

b) What interval centered at a score of 100 contains 50% of the population?

c) For a population of 2,500, how many are expected to have a score above 125?
1
Expert's answer
2020-11-05T17:31:43-0500

a. "P(95<X<110)"

"P(95<X<110)=P(\frac{95-100}{15}<Z<\frac{110-100}{15})"

"=P(Z<0.6667)-P(Z< -0.3333)"

"=0.7475-0.3694=0.3781"

Thus, 37.81% of the population would obtain scores between 95 and 110

b"P(<X)=0.5-0.25" to get the left value

"P(\frac{X-100}{15})=0.25" X is less than 100

"P(Z<-y)=0.25"

"y=\frac{X-100}{15}" from =NORM.S.INV(0.25) excel formula "y=-0.67448975"

thus "-0.67448975=\frac{X-100}{15}"

"X=(-0.67448975\times15)+100=89.88265"

The right limit is "(100-89.88265)+100=110.1173"

the interval {89.88,110.12} contains 50% of the population and is centered at a score of 100

c. P(>125)

"P(\frac{125-100}{15})=P(Z>)=1.667=0.0478"

"2500\times0.0478=119.5"

120 people out of 2500 will score above 125